Four less than six times a certain number is equal to 40 more than one-half of the number. Find the number
Answer by ewatrrr(24785)


Hi,
Let x represent the number
Question states***Write as we Read
6x - 4 = x/2 + 40 |multiplying each term by 2
12x - 8 = x + 80
11x = 88
x = 8
CHECKING our Answer****
48-4 = 4 + 40 = 44